The height of a triangle is 18.6 meters. Its area is 46.5 square meters. Find the length of the triangles base
Answer by cleomenius(959)

area = 1/2 b*h
46.5 m^2 = 1/2 b*18.6
93 = b*18.6
5 m = Base.
1/2 *5 m* (18.6)m = 46.5 m^2.
This does check,
Cleomenius